# IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators

### v3

@@ I Introduction
In an era dominated by online media, understanding and mitigating bias within digital content [1, 2, 3] have become paramount. The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume. In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing media bias, specifically political bias, in online media content, which only be done by journalism experts in the past.
At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles. Leveraging advanced techniques from large language models (LLMs), vector database [4] and IndiVec technique [5], our system employs a two-stage approach to facilitate bias analysis. In the offline stage, IndiTag constructs a bias indicator vector database using LLMs and meticulously designed prompts. These prompts serve as guides for LLMs in generating fine-grained bias indicators, which are then rigorously verified to ensure quality and reliability.
@@
Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias [6], BASIL [7], BABE [8], MFC [9]). These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
@@ II Related Work
Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information [10]. While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases [3], linguistic biases [8], text-level context biases [11], gender biases [12], and racial biases [13]. Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines [14, 1, 15] to more advanced techniques such as Recurrent Neural Networks [2] and pretrained language model-based methods like BERT and RoBERTa [3, 7].
Recent research has explored diverse aspects of media bias, including defining it [16, 17, 18], detecting bias in content [19, 20, 21, 22, 23], understanding bias perception [24, 25, 26, 27], and promoting bias awareness through educational interventions [28, 29, 30, 31]. Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations [32]. This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
@@ III-B Workflow Overview
  • Offline Bias Indicator Database. The vector database is constructed offline to serve as the core module supporting online indicator matching, consisting of sentence embeddings representing bias indicators. Leveraging LLMs, we employ effective prompts to automatically extract bias indicators from a dataset containing bias and corresponding labels (e.g., political left-leaning or right-leaning). Our offline database construction methodology closely follows our previous research paper [5], with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance. Following this extraction process, we store the embeddings of the entire indicator set in the database, ready for use in online bias analysis queries.
@@ III-C Implementation Details
    Designing Prompts for Indicator Generation. We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (see details in our prior work [5]). These prompts cover multiple aspects of media bias assessment, including tone and language, sources and citations, coverage and balance, agenda and framing, and bias in examples and analogies. The detailed prompt is shown below:
@@
    where DESC&EX represents the description and examples of indicator categories, consistent with those used in [5]. Building upon the prompt used for generating the indicators, we design a similar prompt for descriptor generation to ensure that the descriptors of the input text share a consistent style with the indicators stored in our database.
@@
    Strict Clustering-Based Indicator Improvement. In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach [33] to further refine the quality of bias indicators generated by LLMs. Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters according to Agglomerative Clustering [34, 35]. The number of clusters is controlled by the hyper-parameter $\alpha$, which dictates the maximum distance for indicators to be clustered together. Subsequently, we select the indicator nearest to the center of each cluster as the representative indicator. This strict clustering process serves two main purposes: 1) it effectively reduces redundancy among indicators, ensuring that only the most distinct and informative indicators are retained, and 2) it balances the distribution of indicators across various clusters, preventing an overabundance of indicators within specific semantic spaces. By employing this method, we enhance the overall quality and diversity of bias indicators, thereby improving the reliability and accuracy of bias analysis within the IndiTag system. The filtered indicator set is denoted as $\mathcal{I}$.
@@ IV-A Datasets
We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance. Our bias indicator vector database’s construction utilizes the FlipBias dataset [6], sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
Additionally, we evaluate our model on three other datasets: BASIL [7], BABE [8], and the Media Frame Corpus (MFC) [9]. These datasets are relabeled as Biased and Non-Biased instances following [36]. FlipBias: Consists of 2,781 events with articles from left, center, and right perspectives, with biased and non-biased labels. BASIL: Contains 100 sets of articles from various sources, annotated for lexical and informational bias at the span level. BABE: Comprises 3,673 sentences from Media Cloud, annotated for bias by expert annotators. MFC: Contains 37,622 articles labeled as “pro”, “neutral”, or “anti”, with “pro” and “anti” tones considered as biased.
@@ IV-B Comparisons
  • Fine-tuned Language Models. For the Finetune model, we fine-tune pretrained language models, specifically BERT [37] and GPT3.5, using the training set of the FlipBias dataset. We then present the test performance results on the test sets of the four datasets. When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model [37] and the AdamW optimizer [38]. This fine-tuning process was facilitated through Hugging Face [39], and we specifically employed the BertForSequenceClassification model.
@@
  • IndiVec. IndiVec [5] represents the state-of-the-art research in political bias prediction tasks. While our IndiTag shares similarities with IndiVec, we distinguish ourselves through the adoption of a rigorous clustering-based approach to refine the quality of generated bias indicators.

### v4

@@ I Introduction
In an era dominated by online media, understanding and mitigating bias within digital content [yu2008classifying, iyyer2014political, liu2022politics] have become paramount. The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume. In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing media bias, specifically political bias, in online media content, which only be done by journalism experts in the past.
At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles. Leveraging advanced techniques from large language models (LLMs), vector database [peng2023embedding] and IndiVec technique [lin2024indivec], our system employs a two-stage approach to facilitate bias analysis. In the offline stage, IndiTag constructs a bias indicator vector database using LLMs and meticulously designed prompts. These prompts serve as guides for LLMs in generating fine-grained bias indicators, which are then rigorously verified to ensure quality and reliability.
@@
Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias [chen2018learning], BASIL [fan2019plain], BABE [spinde2022neural], MFC [card2015media]). These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
@@ II Related Work
Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information [golbeck2017large]. While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases [liu2022politics], linguistic biases [spinde2022neural], text-level context biases [farber2020multidimensional], gender biases [grosz2020automatic], and racial biases [barikeri2021redditbias]. Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines [evans2007recounting, yu2008classifying, sapiro2019examining] to more advanced techniques such as Recurrent Neural Networks [iyyer2014political] and pretrained language model-based methods like BERT and RoBERTa [liu2022politics, fan2019plain].
Recent research has explored diverse aspects of media bias, including defining it [nakov_survey_2021, spinde_interdisciplinary_2021, spinde2025leveraging], detecting bias in content [liu_politics_2022, lee_neus_2022, fagni_fine-grained_2022, Spinde2021f, lin2023data], understanding bias perception [goldman_friendly_2011, knobloch-westerwick_looking_2009, barnidge_politically_2020, gearhart_hostile_2020], and promoting bias awareness through educational interventions [weeks_incidental_2017, correll_affirmed_2004, kim_hostile_2019, pennycook_psychology_2021]. Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations [lin2025counterspeech]. This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
@@ III-B Workflow Overview
  • Offline Bias Indicator Database. The vector database is constructed offline to serve as the core module supporting online indicator matching, consisting of sentence embeddings representing bias indicators. Leveraging LLMs, we employ effective prompts to automatically extract bias indicators from a dataset containing bias and corresponding labels (e.g., political left-leaning or right-leaning). Our offline database construction methodology closely follows our previous research paper [lin2024indivec], with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance. Following this extraction process, we store the embeddings of the entire indicator set in the database, ready for use in online bias analysis queries.
@@ III-C Implementation Details
    Designing Prompts for Indicator Generation. We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (see details in our prior work [lin2024indivec]). These prompts cover multiple aspects of media bias assessment, including tone and language, sources and citations, coverage and balance, agenda and framing, and bias in examples and analogies. The detailed prompt is shown below:
@@
    where DESC&EX represents the description and examples of indicator categories, consistent with those used in [lin2024indivec]. Building upon the prompt used for generating the indicators, we design a similar prompt for descriptor generation to ensure that the descriptors of the input text share a consistent style with the indicators stored in our database.
@@
    Strict Clustering-Based Indicator Improvement. In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach [feng2025graph] to further refine the quality of bias indicators generated by LLMs. Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters according to Agglomerative Clustering [day1984efficient, feng2024llmedgerefine]. The number of clusters is controlled by the hyper-parameter $\alpha$, which dictates the maximum distance for indicators to be clustered together. Subsequently, we select the indicator nearest to the center of each cluster as the representative indicator. This strict clustering process serves two main purposes: 1) it effectively reduces redundancy among indicators, ensuring that only the most distinct and informative indicators are retained, and 2) it balances the distribution of indicators across various clusters, preventing an overabundance of indicators within specific semantic spaces. By employing this method, we enhance the overall quality and diversity of bias indicators, thereby improving the reliability and accuracy of bias analysis within the IndiTag system. The filtered indicator set is denoted as $\mathcal{I}$.
@@ IV-A Datasets
We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance. Our bias indicator vector database’s construction utilizes the FlipBias dataset [chen2018learning], sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
Additionally, we evaluate our model on three other datasets: BASIL [fan2019plain], BABE [spinde2022neural], and the Media Frame Corpus (MFC) [card2015media]. These datasets are relabeled as Biased and Non-Biased instances following [wessel2023introducing]. FlipBias: Consists of 2,781 events with articles from left, center, and right perspectives, with biased and non-biased labels. BASIL: Contains 100 sets of articles from various sources, annotated for lexical and informational bias at the span level. BABE: Comprises 3,673 sentences from Media Cloud, annotated for bias by expert annotators. MFC: Contains 37,622 articles labeled as “pro”, “neutral”, or “anti”, with “pro” and “anti” tones considered as biased.
@@ IV-B Comparisons
  • Fine-tuned Language Models. For the Finetune model, we fine-tune pretrained language models, specifically BERT [devlin2018bert] and GPT3.5, using the training set of the FlipBias dataset. We then present the test performance results on the test sets of the four datasets. When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model [devlin2018bert] and the AdamW optimizer [loshchilov2017decoupled]. This fine-tuning process was facilitated through Hugging Face [wolf2020transformers], and we specifically employed the BertForSequenceClassification model.
@@
  • IndiVec. IndiVec [lin2024indivec] represents the state-of-the-art research in political bias prediction tasks. While our IndiTag shares similarities with IndiVec, we distinguish ourselves through the adoption of a rigorous clustering-based approach to refine the quality of generated bias indicators.